OpenStudy (anonymous):

f(x) = x^3 - ax^2 - bx - 3a and g(x) = x^3 -(a-2)x^2 - bx -3b have same quadratic factor. How to find a and b value?

ganeshie8 (ganeshie8):

since they have a common quadratic factor, there will be two common roots as well : say, \(p, q, r\) are roots of \(f(x)\) and \(p,q,s\) are roots of \(g(x)\) \(f(x) = x^3 - ax^2 - bx - 3a = (x-p)(x-q)(x-r)\) \(g(x) = x^3 -(a-2)x^2 - bx -3b =(x-p)(x-q)(x-s)\)

ganeshie8 (ganeshie8):

equation \(x\) coefficient in both equations, gives :- \(-b = pq + qr + rp\) \(-b = pq + qs + sp\)

ganeshie8 (ganeshie8):

subtracting both :- \(0 = q(r-s)+ p(r-s) = (r-s)(q+p)\) for this to be true, atleast one factor must equal 0 \(r=s, \) or \(p = -q\)

OpenStudy (ranga):

If r = s, f(x) and g(x) will be one and the same function. So p = -q.

ganeshie8 (ganeshie8):

Oh ya, so we can discard r=s and pursue p = -q case

ganeshie8 (ganeshie8):

still its not fully done... u sure u can work the rest ha ? :)

ganeshie8 (ganeshie8):

i see lot of algebra still.... good luck !

OpenStudy (anonymous):

this is my lil' sister homework, she is in high school, but i don't know why its so difficult :(

ganeshie8 (ganeshie8):

she must be in12th grade ?

OpenStudy (ranga):

There are 6 unknowns: a, b, p, q, r, s and there are 6 equations, equating the coefficients of x^2, x and constant in to the two functions coefficients.

OpenStudy (anonymous):

I'm stuck -__-